# Incidence estimates for tubes in complex space

### v2

@@ Abstract.
In this paper, we prove a complex version of the incidence estimate of Guth, Solomon and Wang [GSW19] for tubes obeying certain strong spacing conditions, and we use one of our new estimates to resolve a discretized variant of Falconer’s distance set problem in $\mathbb{C}^{2}$.
@@ 1. Introduction
  • the Kakeya problem in dimension three under the Wolff axioms is true over $\mathbb{R}$ [WZ25] but false over $\mathbb{C}$ because of the Heisenberg example [KŁT19];
  • the Falconer distance problem and its cousins, the Erdös ring problem and the Furstenberg problem, have counterexamples over $\mathbb{C}$ because of the existence of the quadratic subfield $\mathbb{R}$ [Wol03], but are still open over $\mathbb{R}$;
  • the Szemerédi-Trotter theorem is true over both $\mathbb{R}$ [ST83] and $\mathbb{C}$ [Tót15];
  • the Erdős distinct distances problem, with an $\varepsilon$-loss, is true over both $\mathbb{R}$ [GK15] and $\mathbb{C}$ [SZ21].
The results in this paper are complex analogues of the incidence estimates and distance estimates obtained in [GSW19], giving some more examples of problems that have positive answers over both $\mathbb{R}$ and $\mathbb{C}$.
@@
Fix $21<s<$ 2. Let $E$ be a set of $\sim\delta^{-s}$ many $\delta$-balls in $B_{\mathbb{C}^{2}}(\boldsymbol{0},1)$ with at most $N$ $\delta$-balls in each ball of radius $\delta^{s/4}$. Then, the number of disjoint $\delta$-balls needed to cover the difference set
@@ 2. Definitions, notations and conventions
[…]
Note that for $2$-planes in $\mathbb{R}^{2n}$ corresponding to complex tubes, the two principal angles are equal [^3] Let’s assume WLOG that $\boldsymbol{t}_{1}=\boldsymbol{t}_{2}=\boldsymbol{0}$. Suppose the first principal angle is attained for the pair of unit vectors $\left(\boldsymbol{v}_{1},\boldsymbol{v}_{2}\right)\in e^{i\mathbb{R}}\boldsymbol{u}_{1}\times e^{i\mathbb{R}}\boldsymbol{u}_{2}$. Since the orthogonal complement of $\iota\left(\boldsymbol{v}_{j}\right)$ in $\iota\left(\ell_{j}\right)$ is $\operatorname{span}_{\mathbb{R}}\left(\iota\left(i\boldsymbol{v}_{j}\right)\right)$ for each $j\in\{1,2\}$, the second principal angle between $\iota\left(\ell_{1}\right)$ and $\iota\left(\ell_{2}\right)$ is $\arccos\left(\iota\left(i\boldsymbol{v}_{1}\right)\cdot\iota\left(i\boldsymbol{v}_{2}\right)\right)=\arccos\left(\iota\left(\boldsymbol{v}_{1}\right)\cdot\iota\left(\boldsymbol{v}_{2}\right)\right).$ Hence the study of incidences between complex lines in $\mathbb{C}^{n}$ is different from the study of incidences between general $2$-planes in $\mathbb{R}^{2n}$.
@@ 3. A bridge between different scales
#{enlarged dual slabs passing through $\#\{\text{enlarged dual slabs passing through }\boldsymbol{\omega}\}$ $\sim\frac{D^{2(n-1)}\cdot\left(\lambda D^{-1}\right)^{2}\|\boldsymbol{\omega}\|^{2(n-2)}}{\|\boldsymbol{\omega}\|^{2(n-1)}}$
@@
#$\#\left\{T\in\mathbb{T}:q\cap\mathcal{N}_{\lambda}(T)\neq\varnothing\right\}\gtrsim E\lambda^{-2}\rho^{-2n}.$
@@
[…]
@@ 4. Proof of the main theorems
Now, assume that the following estimate holds when either $\widetilde{r}\geqslant 2r$ or $\widetilde{\delta}\geqslant\delta^{-\varepsilon^{10}}\delta$ or both, and $\widetilde{r}\geqslant c_{1}(\varepsilon)\,\widetilde{\delta}^{2(n-1)-\varepsilon}\left|\widetilde{\mathbb{T}}\right|$:
@@
$I(10\widetilde{\mathbb{T}}_{M_{0},\max},\widetilde{P})\supseteq I(\widetilde{\mathbb{T}}_{M_{0}},\widetilde{P}),$
@@
$P_{r,\alpha}(\mathbb{T}):=\left\{p\in P_{r}(\mathbb{T}):\begin{subarray}{c}\text{The maximal angle between complex tubes}\\ \text{of $\mathbb{T}$ passing through the $\delta$-ball $p$ is $\sim\alpha$.}\end{subarray}\right\}$
@@
    For this reason, we may as well assume $W\leqslant\delta^{-\frac{1}{2}+\frac{\varepsilon}{8}}$ in the rest of part (b). Then, $W^{-2}\geqslant\delta^{-\frac{\varepsilon}{4}+1}\gg\delta$. Let $\widetilde{\mathbb{T}}$ be the collection of complex $W^{-2}$-tubes obtained by thickening each complex $\delta$-tube in $\mathbb{T}$, and let $\widetilde{\mathbb{T}}_{\max}\subseteq\widetilde{\mathbb{T}}$ be a maximal subcollection of essentially distinct $W^{-2}$-tubes. Since $W^{-2}<W^{-1}$, the spacing condition of $\mathbb{T}$ implies that $\left|\widetilde{\mathbb{T}}_{\max}\right|\geqslant N_{0}^{-1}\left|\widetilde{\mathbb{T}}\right|$. Let $Q$ be a minimal covering of $B_{\mathbb{C}^{n}}(\boldsymbol{0},1)$ consisting of $W^{-2}$-balls; and for dyadic $X,M\geqslant 1$, let $Q_{X,M}$ be the collection of $W^{-2}$-balls in $Q$ that contains $\sim X$ of the $\delta$-balls in $P_{r,\alpha}(\mathbb{T})$ and intersects $\sim M$ of [^8] At least $M$ and strictly less than $2M$ of the complex tubes, to be precise. the complex $\delta$-tubes in $\mathbb{T}$. We shall note that the set $Q_{X,M}$ is nonempty only if $X\lesssim\left(\frac{W^{-2}}{\delta}\right)^{\!2n}$ and $r\leqslant M\lesssim W^{4(n-1)}$. Then, by dyadic pigeonholing twice, we can find some particular $X_{0}$ and $M_{0}$ such that $\bigcup_{q\in Q_{X_{0},M_{0}}}q$ covers a $\gtrsim_{n}(-\log\delta)^{-2}$ fraction of the $\delta$-balls in $P_{r,\alpha}(\mathbb{T})$. By definition, each $\delta$-ball in $P_{r,\alpha}(\mathbb{T})$ is intersected by two complex tubes in $\mathbb{T}$ with directions differing by $\sim\alpha$. Hence, the number of $\delta$-balls from $P_{r,\alpha}(\mathbb{T})$ contained in each $q$ is $\lesssim\alpha^{-2}M_{0}^{2}$; and consequently we have
@@
    with the second inequality following from (4). When $-2+\frac{n+1}{n-1}\geqslant 0$, i.e. when $n=2\text{ or }3$,
@@
    $\widetilde{P}:=\bigcup_{\mathcal{q}\text{ thick}}\widetilde{P}_{\mathcal{q}}.$
    Then, each $\widetilde{p}\in\widetilde{P}$ intersects $\gtrsim N_{0}E_{0}D^{\frac{n-1}{50n}\varepsilon-2n\varepsilon^{3}}$ of the complex tubes $T\in\mathbb{T}$. To apply the induction hypothesis, we fatten each complex $\delta$-tube $T\in\mathbb{T}$ into a complex tube $\widetilde{T}$ of width $\sim D^{\frac{\varepsilon}{100n}}\delta$, and call the collection of all these complex tubes $\widetilde{\mathbb{T}}$. Then, $\widetilde{P}\subseteq P_{\widetilde{r}}(\widetilde{\mathbb{T}})$ for $\widetilde{r}$ a sufficiently small constant fraction of $H_{0}E_{0}D^{\frac{n-1}{50n}\varepsilon-2n\varepsilon^{3}}$, and for such $\widetilde{r}$ we have
@@
    $\sum_{\mathcal{q}\text{ thick}}\left|\widetilde{P}_{\mathcal{q}}\right|\leqslant\left|P_{\widetilde{r}}\left(\widetilde{\mathbb{T}}\right)\right|$ $\lesssim\widetilde{\delta}^{-\varepsilon}\widetilde{r}^{-\frac{n+1}{n-1}}\left|\widetilde{\mathbb{T}}\right|^{\frac{n}{n-1}}$
@@ 5. Application to a variant of Falconer’s distance set problem
Fix $21<s<$ 2. Let $E$ be a collection of $\sim N\delta^{-s}$ many101010say $\frac{1}{2}\left|B_{\mathbb{C}^{2}}(\vec{0},1)\right|N\delta^{-s}<|E|<2\left|B_{\mathbb{C}^{2}}(\vec{0},1)\right|N\delta^{-s}$ $\delta$-balls in $B_{\mathbb{C}^{2}}(\vec{0},1)$ with at most $N$ $\delta$-balls in each ball of radius $\delta^{s/4}$. How many finitely overlapping $\delta$-balls are needed to cover the difference set
@@
The intuition behind this definition comes from the Elekes-Sharir framework used in the distinct distance problem in $\mathbb{R}^{2}$ (see e.g. [GK15]): Over $\mathbb{R}$ instead of $\mathbb{C}$, the orthogonal projection of $l_{\vec{p}_{1},\vec{p}_{2}}$ onto the first two coordinates is the perpendicular bisector of the segment between the points $\vec{p}_{1}$ and $\vec{p}_{2}$ in $\mathbb{R}^{2}$, and the line $l_{\vec{p}_{1},\vec{p}_{2}}$ itself parametrizes all the rigid motions mapping $\vec{p}_{1}$ to $\vec{p}_{2}$. Two things to note:
@@
$\begin{dcases}\frac{\left(x_{\vec{\sigma}_{1}}+r_{1}\right)+\left(x_{\vec{\sigma}_{2}}+r_{2}\right)}{2}+(z+c)\frac{\left(y_{\vec{\sigma}_{1}}+s_{1}\right)+\left(y_{\vec{\sigma}_{2}}+s_{2}\right)}{2}&\\ &=\frac{x_{\vec{\sigma}_{1}}+x_{\vec{\sigma}_{2}}}{2}+z\frac{y_{\vec{\sigma}_{1}}-y_{\vec{\sigma}_{2}}}{2}+a\\ \frac{\left(y_{\vec{\sigma}_{1}}+s_{1}\right)+\left(y_{\vec{\sigma}_{2}}+s_{2}\right)}{2}+(z+c)\frac{\left(x_{\vec{\sigma}_{1}}+r_{1}\right)+\left(x_{\vec{\sigma}_{2}}+r_{2}\right)}{2}&\\ &=\frac{x_{\vec{\sigma}_{1}}+x_{\vec{\sigma}_{2}}}{2}+z\frac{y_{\vec{\sigma}_{1}}-y_{\vec{\sigma}_{2}}}{2}+b\end{dcases}.$
@@
Here, we set $W\sim\delta^{-s/4}$ (rather than $W=\delta^{-s/4}$) to account for various implied constants that appear when proving that a spacing condition for the $\delta$-balls results in a spacing condition for the associated tubes. Ultimately, we want to make $W^{-1}$ a small positive multiple of $\delta^{s/4}$ (e.g. by $1/1000$); thus we should think of $W$ as a multiple of $\delta^{-s/4}$ by a constant $>1$. One could alternatively read the proposition as the statement, “there exists $C>1$ so that if $W=C\delta^{-s/4}$ and $N_{0}=N^{2}$, then the collection $\mathbb{T}$ meets the spacing condition of Corollary 4.3.”
@@
#$\#\Delta(E^{\prime},E^{\prime\prime})\geqslant\frac{(\#E^{\prime}\cdot\#E^{\prime\prime})^{2}}{|Q|}\gtrsim|\log\delta|^{-1}\delta^{-s+\varepsilon}.$
@@
0.30.3 0.30.3 $\geqslant\left\|\frac{\vec{p}_{1}-\vec{p}_{3}}{2}-z\frac{\vec{p}_{2}-\vec{p}_{4}}{2}\right\|$
@@
To summarize, in order for $\theta_{z}$ to be small, the first two coordinates of $\boldsymbol{v}_{\vec{p}_{1},\vec{p}_{3}}$ and of $\boldsymbol{v}_{\vec{p}_{2},\vec{p}_{4}}$ force $z$ to be closer to $-1$ while the last coordinate of $\boldsymbol{v}_{\vec{p}_{1},\vec{p}_{3}}$ and of $\boldsymbol{v}_{\vec{p}_{2},\vec{p}_{4}}$ force $z$ to be closer to $1$; yet the two cannot be achieved at the same time; thus $\theta_{z}$ cannot be too small. In conclusion,
@@
The distance between $l_{\vec{p}_{1},\vec{p}_{3}}$ and $l_{\vec{p}_{2},\vec{p}_{4}}$ is $<\delta/2$. Consequently, the complex tubes $T_{q_{1},q_{3}}$ and $T_{q_{2},q_{4}}$ intersect in (essentially) a $\delta$-ball in $B_{\mathbb{C}_{n}}\left(\boldsymbol{0},1\right)$
@@
We can think of partitioning $E$ as assigning a color to each $\delta$-ball in $E$. We adapt a graph-theoretic argument that we first encountered in [FR24], in the proof of their Proposition 4.1. We build a graph whose vertex set is the set of centres of the $\delta$-balls in $E$ and place edges so that two vertices have an edge between them if and only if the corresponding $\delta$-balls are both contained in a larger $\delta^{s/4}$-ball. The resulting graph has maximum degree $N-1$, so by Brooks’ Theorem, the chromatic number of the graph is at most $N$; that is, we can color the $\delta$-balls of $E$ using $N$ colors in a way that guarantees that no two $\delta$-balls of the same color are in any $\delta^{s/4}$-ball.
@@
By 5.12 — which is stated and proved below — the perpendicular distance from $\boldsymbol{G}_{1,3}(\alpha,\beta)$ to the plane $P_{2,4}$ is comparable to the distance from $\boldsymbol{G}_{1,3}(\alpha,\beta)$ to $\boldsymbol{G}_{2,4}(\alpha,\beta)$; specifically,
@@
$\begin{split}4(\delta^{s/4})^{2}&\leqslant\left(\frac{1}{8}+1\right)\left\|(\vec{\sigma}_{1}-\vec{\sigma}_{3})-(\vec{\sigma}_{2}-\vec{\sigma}_{4})\right\|^{2}=\frac{9}{8}^{\,}\left\|(\vec{\sigma}_{1}-\vec{\sigma}_{3})-(\vec{\sigma}_{2}-\vec{\sigma}_{4})\right\|^{2}.\\ \end{split}$

### v3

@@ Abstract.
In this paper, we prove a complex version of the incidence estimate of Guth, Solomon and Wang [WellSpacedTubes] for tubes obeying certain strong spacing conditions, and we use one of our new estimates to resolve a discretized variant of Falconer’s distance set problem in $\mathbb{C}^{2}$.
@@ 1. Introduction
  • the Kakeya problem in dimension three under the Wolff axioms is true over $\mathbb{R}$ [Kakeya_3dim_WolffAxioms_Real_new] but false over $\mathbb{C}$ because of the Heisenberg example [Kakeya_3dim_WolffAxioms_Old];
  • the Falconer distance problem and its cousins, the Erdös ring problem and the Furstenberg problem, have counterexamples over $\mathbb{C}$ because of the existence of the quadratic subfield $\mathbb{R}$ [WolffRecentKakeya], but are still open over $\mathbb{R}$;
  • the Szemerédi-Trotter theorem is true over both $\mathbb{R}$ [SzT_Original] and $\mathbb{C}$ [SzT_Complex];
  • the Erdős distinct distances problem, with an $\varepsilon$-loss, is true over both $\mathbb{R}$ [ErdosDistinctDist_Real] and $\mathbb{C}$ [ErdosDistinctDist_Complex].
The results in this paper are complex analogues of the incidence estimates and distance estimates obtained in [WellSpacedTubes], giving some more examples of problems that have positive answers over both $\mathbb{R}$ and $\mathbb{C}$.
@@
Fix $2<s<4$. Let $E$ be a set of $\sim\delta^{-s}$ many $\delta$-balls in $B_{\mathbb{C}^{2}}(\boldsymbol{0},1)$ with at most NN $\delta$-balls in each ball of radius $\delta^{s/4}$. Then, the number of disjoint $\delta$-balls needed to cover the difference set
@@ 2. Definitions, notations and conventions
    Define the angle numerically to be $\arccos|\langle\boldsymbol{u}_{1},\boldsymbol{u}_{2}\rangle|$, where $\langle\cdot,\cdot\rangle$ is the usual inner product on $\mathbb{C}^{n}$; this is also the Fubini-Study distance, up to a scaling constant, between the equivalent classes of $\boldsymbol{u}_{1}$ and $\boldsymbol{u}_{2}$ in $\mathbb{C}\mathbb{P}^{n-1}$.
Note that for $2$-planes in $\mathbb{R}^{2n}$ corresponding to complex tubes, the two principal angles are equal [^3] Let’s assume WLOG that $\boldsymbol{t}_{1}=\boldsymbol{t}_{2}=\boldsymbol{0}$. Suppose the first principal angle is attained for the pair of unit vectors $\left(\boldsymbol{v}_{1},\boldsymbol{v}_{2}\right)\in e^{i\mathbb{R}}\boldsymbol{u}_{1}\times e^{i\mathbb{R}}\boldsymbol{u}_{2}$. Since the orthogonal complement of $\iota\left(\boldsymbol{v}_{j}\right)$ in $\iota\left(\ell_{j}\right)$ is $\operatorname{span}_{\mathbb{R}}\left(\iota\left(i\boldsymbol{v}_{j}\right)\right)$ for each $j\in\{1,2\}$, the second principal angle between $\iota\left(\ell_{1}\right)$ and $\iota\left(\ell_{2}\right)$ is $\arccos\left(\iota\left(i\boldsymbol{v}_{1}\right)\cdot\iota\left(i\boldsymbol{v}_{2}\right)\right)=\arccos\left(\iota\left(\boldsymbol{v}_{1}\right)\cdot\iota\left(\boldsymbol{v}_{2}\right)\right).$ . Hence the study of incidences between complex lines in $\mathbb{C}^{n}$ is different from the study of incidences between general $2$-planes in $\mathbb{R}^{2n}$.
@@ 3. A bridge between different scales
$\#\{\text{enlarged dual slabs passing through }\boldsymbol{\omega}\}$ $\sim\frac{D^{2(n-1)}\cdot\left(\lambda D^{-1}\right)^{2}\|\boldsymbol{\omega}\|^{2(n-2)}}{\|\boldsymbol{\omega}\|^{2(n-1)}}$
@@
$\left\{T\in\mathbb{T}:q\cap\mathcal{N}_{\lambda}(T)\neq\varnothing\right\}\gtrsim E\lambda^{-2}\rho^{-2n}.$
@@
$\sim\int(1-\eta(\boldsymbol{\omega}))\left|\sum_{\theta:\lambda T_{\theta}^{*}\ni\boldsymbol{\omega}}\widehat{g_{\theta}}(\boldsymbol{\omega})\right|^{2}d\boldsymbol{\omega}+\int(1-\eta(\boldsymbol{\omega}))\left|\sum_{\theta:\lambda T_{\theta}^{*}\not\ni\boldsymbol{\omega}}\widehat{g_{\theta}}(\boldsymbol{\omega})\right|^{2}d\boldsymbol{\omega}$
@@ 4. Proof of the main theorems
[…]
@@
[…]
@@
[…]
@@
[…]
@@
    with the second inequality following from (4). When $2+\frac{n+1}{n-1}\geqslant 0$, i.e. when $n=2\text{ or }3$,
@@
[…]
@@
[…]
@@ 5. Application to a variant of Falconer’s distance set problem
Fix $2<s<4$. Let $E$ be a collection of $\sim N\delta^{-s}$ many101010say $\frac{1}{2}\left|B_{\mathbb{C}^{2}}(\vec{0},1)\right|N\delta^{-s}<|E|<2\left|B_{\mathbb{C}^{2}}(\vec{0},1)\right|N\delta^{-s}$ $\delta$-balls in $B_{\mathbb{C}^{2}}(\vec{0},1)$ with at most NN $\delta$-balls in each ball of radius $\delta^{s/4}$. How many finitely overlapping $\delta$-balls are needed to cover the difference set
@@
The intuition behind this definition comes from the Elekes-Sharir framework used in the distinct distance problem in $\mathbb{R}^{2}$ (see e.g. [ErdosDistinctDist_Real]): Over $\mathbb{R}$ instead of $\mathbb{C}$, the orthogonal projection of $l_{\vec{p}_{1},\vec{p}_{2}}$ onto the first two coordinates is the perpendicular bisector of the segment between the points $\vec{p}_{1}$ and $\vec{p}_{2}$ in $\mathbb{R}^{2}$, and the line $l_{\vec{p}_{1},\vec{p}_{2}}$ itself parametrizes all the rigid motions mapping $\vec{p}_{1}$ to $\vec{p}_{2}$. Two things to note:
@@
$\begin{dcases}\frac{\left(x_{\vec{\sigma}_{1}}+r_{1}\right)+\left(x_{\vec{\sigma}_{2}}+r_{2}\right)}{2}+(z+c)\frac{\left(y_{\vec{\sigma}_{1}}+s_{1}\right)+\left(y_{\vec{\sigma}_{2}}+s_{2}\right)}{2}&\\ &\hskip-80.00012pt=\frac{x_{\vec{\sigma}_{1}}+x_{\vec{\sigma}_{2}}}{2}+z\frac{y_{\vec{\sigma}_{1}}-y_{\vec{\sigma}_{2}}}{2}+a\\ \frac{\left(y_{\vec{\sigma}_{1}}+s_{1}\right)+\left(y_{\vec{\sigma}_{2}}+s_{2}\right)}{2}+(z+c)\frac{\left(x_{\vec{\sigma}_{1}}+r_{1}\right)+\left(x_{\vec{\sigma}_{2}}+r_{2}\right)}{2}&\\ &\hskip-80.00012pt=\frac{x_{\vec{\sigma}_{1}}+x_{\vec{\sigma}_{2}}}{2}+z\frac{y_{\vec{\sigma}_{1}}-y_{\vec{\sigma}_{2}}}{2}+b\end{dcases}.$
@@
Here, we set $W\sim\delta^{-s/4}$ (rather than $W=\delta^{-s/4}$) to account for various implied constants that appear when proving that a spacing condition for the $\delta$-balls results in a spacing condition for the associated tubes. Ultimately, we want to make $W^{-1}$ a small positive multiple of $\delta^{s/4}$ (e.g. by $1/1000$); thus we should think of $W$ as a multiple of $\delta^{-s/4}$ by a constant >1>1. One could alternatively read the proposition as the statement, “there exists $C>1$ so that if $W=C\delta^{-s/4}$ and $N_{0}=N^{2}$, then the collection $\mathbb{T}$ meets the spacing condition of Corollary 4.3.”
@@
$\Delta(E^{\prime},E^{\prime\prime})\geqslant\frac{(\#E^{\prime}\cdot\#E^{\prime\prime})^{2}}{|Q|}\gtrsim|\log\delta|^{-1}\delta^{-s+\varepsilon}\geqslant\delta^{-2+\varepsilon}.$
@@
$0.3$ $\geqslant\left\|\frac{\vec{p}_{1}-\vec{p}_{3}}{2}-z\frac{\vec{p}_{2}-\vec{p}_{4}}{2}\right\|$
@@
To summarize, in order for $\theta_{z}$ to be small, the first two coordinates of $\boldsymbol{v}_{\vec{p}_{1},\vec{p}_{3}}$ and of $\boldsymbol{v}_{\vec{p}_{2},\vec{p}_{4}}$ force $z$ to be closer to 1-1 while the last coordinate of $\boldsymbol{v}_{\vec{p}_{1},\vec{p}_{3}}$ and of $\boldsymbol{v}_{\vec{p}_{2},\vec{p}_{4}}$ force $z$ to be closer to $1$; yet the two cannot be achieved at the same time; thus $\theta_{z}$ cannot be too small. In conclusion,
@@
The distance between $l_{\vec{p}_{1},\vec{p}_{3}}$ and $l_{\vec{p}_{2},\vec{p}_{4}}$ is $\delta/2$. Consequently, the complex tubes $T_{q_{1},q_{3}}$ and $T_{q_{2},q_{4}}$ intersect in (essentially) a $\delta$-ball in $B_{\mathbb{C}_{n}}\left(\boldsymbol{0},1\right)$
@@
We can think of partitioning $E$ as assigning a color to each $\delta$-ball in $E$. We adapt a graph-theoretic argument that we first encountered in [Fu_Ren_Incidences], in the proof of their Proposition 4.1. We build a graph whose vertex set is the set of centres of the $\delta$-balls in $E$ and place edges so that two vertices have an edge between them if and only if the corresponding $\delta$-balls are both contained in a larger $\delta^{s/4}$-ball. The resulting graph has maximum degree N1N-1, so by Brooks’ Theorem, the chromatic number of the graph is at most $N$; that is, we can color the $\delta$-balls of $E$ using $N$ colors in a way that guarantees that no two $\delta$-balls of the same color are in any $\delta^{s/4}$-ball.
@@
By 5.13 — which is stated and proved below — the perpendicular distance from $\boldsymbol{G}_{1,3}(\alpha,\beta)$ to the plane $P_{2,4}$ is comparable to the distance from $\boldsymbol{G}_{1,3}(\alpha,\beta)$ to $\boldsymbol{G}_{2,4}(\alpha,\beta)$; specifically,
@@
[…]